### v2

# Enhancing the hyperpolarizability of crystals with quantum geometry

Abstract
We demonstrate that higher-order electric susceptibilities in crystals can be enhanced and understood through nontrivial topological invariants and quantum geometry, using one-dimensional $\pi$-conjugated chains as representative model systems. First, we show that the crystalline-symmetry-protected topology of these chains imposes a lower bound on their quantum metric and hyperpolarizabilities. Second, we employ numerical simulations to reveal the tunability of nonlinear, quantum geometry-driven optical responses in various one-dimensional crystals in which band topology can be externally controlled. Third, we develop a semiclassical picture to deliver an intuitive understanding of these effects. Our findings offer a firm interpretation of otherwise elusive experimental observations of colossal hyperpolarizabilities and establish guidelines for designing topological materials of any dimensionality with enhanced nonlinear optical properties.
Introduction.— Nonlinear optics offers a rich platform to explore a broad range of intriguing phenomena, with implications ranging from fundamental science to engineering Boyd (2019). One-dimensional crystals, most notably $\pi$-conjugated linear chains Agrawal et al. (1978); Dalton et al. (1987); Bredas et al. (1994), constitute prime examples to probe enhanced optical responses at nonlinear orders. Because these crystals are centrosymmetric, i.e., they feature an inversion symmetry, their optical responses occur at odd orders in optical fields. The first- and third-order optical responses of one-dimensional $\pi$-conjugated chains have long been investigated, both experimentally  Fincher et al. (1979); Beratan et al. (1987); Sinclair et al. (1988); Samuel et al. (1994); Ledoux et al. (1999); Bredas et al. (1994) and computationally Hermann and Ducuing (1974); Rustagi and Ducuing (1974); Soos and Hayden (1989); Villesuzanne et al. (1992); Huang et al. (1992); Shuai and Brédas (1992, 1992); Schmidt and Springborg (1999); Hu et al. (2007); Lacivita et al. (2012); Luzanov (2023). Of central interest in this context is the third-order hyperpolarizability, i.e., the third-order electric susceptibility $\chi^{(3)}$ in the static limit of vanishing optical field frequency, $\omega\rightarrow 0$, which quantifies the nonlinear polarization response, ${P}^{(3)}=\chi^{(3)}E_{x}^{3}$, to electric field $E_{x}$ Pronin et al. (1995). In $\pi$-conjugated chains, $\chi^{(3)}$ is known to acquire colossal values. For example, $\chi^{(3)}$ was inferred to be as high as $10^{-18}$ m2/V2 in polyacetylene Boyd (2019), $10^{-19}$ m2/V2 in polydiacetylene, and $10^{-20}$ m2/V2 in polyarylenes Bredas et al. (1994). This unusually large hyperpolarizability has been heuristically ascribed to the responses of $p_{z}$-electrons forming the conjugated $\pi$-bonds Boyd (2019). However, a general theoretical understanding of the phenomenological origin of these nonlinear optical properties remains yet to be achieved Yu et al. (1989a); Hagler and Heeger (1994); Takahashi (1995); Yaron (1996); Takahashi (1997a, b); Matsuzaki et al. (1998); Margulis and Gaiduk (1998); Jiang and Xu (2005); Xu and Jiang (2005); Díaz-Ponce (2017).
Figure 1: Crystalline topology and quantum geometrically-enhanced hyperpolarizability in one-dimensional crystals. (a) Schematic illustration of a one-dimensional inversion-symmetric model system with a topological invariant $\mathbb{Z}_{2}$ reflected by a Wannier charge center shift $\langle x\rangle$ and by the presence of a symmetry-protected exponentially-localized edge state, |$\ket{\psi_{\text{edge}}}$. The model system with unit cell parameter $a$ realizes a bipartite basis $A,B$, and effective hoppings $t_{1},t_{2}$ across bonds with lengths $a_{1},a_{2}$, and is coupled to an optical field $E_{x}(\omega)$ introduced by photons with energy $\hbar\omega$. The minimal spread of the Wannier function $\langle\Delta x^{2}\rangle$ is topologically-bounded, and evaluates to the band-averaged quantum metric $\langle g_{xx}\rangle$ Marzari and Vanderbilt (1997), which determines the nonlinear polarizability. (b) Control of the band gap and band topology with lattice strain ($\lambda$) across a topological phase transition (TPT). (c) Discontinuous control of the nonlinear optical properties, such as the time-dependent third-order polarization $P^{(3)}(t)$ from a three-photon coupling, across a strain-induced TPT.
Figure 2: Quantum geometry and (non-)linear optics in one-dimensional crystals. (a) Quantum metric (red) over momentum-space as a function of the lattice strain parameter $(\lambda)$ realized in the effective low-energy Hamiltonian, Eq. (5), with hoppings $t_{1}(0)=2.15~{}\text{eV}$, $t_{2}(0)=2.85~{}\text{eV}$ [$t_{2}(0)/t_{1}(0)=1.33$]. The critical strain $(\lambda=\lambda_{c})$ induces a topological phase transition across which the metric scales up at the band edge at $k=\pm\pi/a(\lambda)$, overcoming the minimal average metric value per BZ (blue) necessary for satisfying the topological bound. (b) Band-averaged metric $\langle g_{xx}\rangle$ as a function of strain and hopping amplitude ratio $t_{2}(0)/t_{1}(0)$. The horizontal line corresponds to the chain with $t_{2}(0)/t_{1}(0)=1.33$. The corresponding optical absorptivity is manifestly larger and topologically lower-bounded in the topologically nontrivial phase. (c) Spectral resolution of strain-dependent optical conductivity $\sigma_{xx}(\omega)$ and cutoff ($\Lambda$) dependent optical weights $W_{xx}(\Lambda)=\int^{\Lambda}_{0}~{}\mathop{}\!\mathrm{d}\omega~{}\frac{\sigma_{xx}(\omega)}{\omega}$ (black), as a function of the bounding value for $\Lambda\rightarrow\infty$ (blue) in a topological chain with $t_{2}/t_{1}=1.33$. (d) Lattice strain control of the topologically-bounded hyperpolarizability $\chi^{(3)}$. Under the relatively-high values of strain $\lambda/\lambda_{0}$, where $\lambda_{0}$ is a unit strain, the anomalous hyperpolarizability of the $\mathbb{Z}_{2}$-topological phase can be quenched. In the subcritical strain limit, we find that due to distinct realized quantum metrics, the $\chi^{(3)}$ of a trivial phase (red) is multiple times smaller than the $\chi^{(3)}$ of a topological phase (blue) with identical band energies. At vanishing strain, we note that the effective model fitted to the $t_{2}/t_{1}$ of polyacetylene yields $\chi^{(3)}=8$$\times$ 10-19 m2/V2, in excellent agreement with the corresponding experimental value, $\chi^{(3)}\approx$ 10-18 m2/V2 Fincher et al. (1979); Dalton et al. (1987); Bredas et al. (1994).
In this Letter, we bridge this gap by deploying the modern notions of band topology and quantum geometry. We show that the crystalline-symmetry-protected topology of $\pi$-conjugated chains induces a lower bound on the quantum metric, and thus on the hyperpolarizability. Then, we use numerical simulations to demonstrate the tunability of nonlinear, quantum geometry-driven optical responses of a variety of one-dimensional crystals in which the band topology is controllable by an external parameter, such as lattice strain. Finally, we propose a general semiclassical picture to explain these remarkable optical properties, based on the scaling of quantum metric and band gaps with external parameters. Importantly, our theory offers a topological explanation for the distinct hyperpolarizabilities of the cis and trans isomers of $\pi$-conjugated linear polymers Sinclair et al. (1988). Unlike the cis configuration, the topological windings of Bloch eigenvector in the trans configuration support the geometric enhancements of the optical matrix elements associated with the nonlinear optical responses. Overall, our findings provide a novel interpretation of otherwise elusive experimental observations, along with guiding principles for identifying and engineering materials for nonlinear optics.
Quantum geometry and topology in one dimension.— Over the past several years, crystalline topologies have been thoroughly investigated Fu (2011); Slager (2019); Slager et al. (2012); Kruthoff et al. (2017); Po et al. (2017); Bradlyn et al. (2017), culminating in a rather uniform view upon comparing momentum space constraints Kruthoff et al. (2017) and real space (Wannierizability) conditions Po et al. (2017); Bradlyn et al. (2017). Within this framework, taking into account the fact that an exponentially-localized Wannier basis Marzari et al. (2012) can be always constructed Kohn (1959) in one spatial dimension, it is directly inferred that the only form of nontrivial electronic topology is that of an obstructed limit. The topology Schnyder et al. (2008); Kitaev et al. (2009) arises from the interplay of band wavefunction (|$\ket{\psi_{n{\bf k}}}=e^{i{\bf k}\cdot{\bf r}}\ket{u_{n{\bf k}}}$) windings with additional symmetries, such as chiral or inversion ($\mathcal{I}$) symmetry Fu and Kane (2007). Under $\mathcal{I}$-symmetry, in nonmagnetic systems, the topological invariant is determined by the band parity eigenvalues, $\delta_{k}=\pm 1$ at inversion-symmetric momenta, $k=0,\pi/a$, where $a$ is the lattice constant, which coincides with the Fu-Kane formula Fu and Kane (2007): ${\nu_{\text{FK}}=(1-\delta_{0}\delta_{\pi/a})/2\in\mathbb{Z}_{2}}$. At the same time, the band topology of obstructed insulators provides for the minimal spread $\langle\Delta x^{2}\rangle$ of the exponentially-localized Wannier functions Marzari and Vanderbilt (1997), as schematically presented in Fig. 1(a). Furthermore, this can be understood with the notions of quantum geometry Provost and Vallee (1980), and can be related to topological bounds Herzog-Arbeitman et al. (2022).
We now elucidate how, in one-dimensional (1D) crystals, the crystalline topology gives rise to quantum geometry and, as a direct consequence, enhances their optical responses. We define a non-Abelian multiband quantum geometric tensor, ${[\textbf{Q}^{(m)}_{ij}]_{np}=\braket{\partial_{k_{i}}u_{n{\bf k}}}{u_{m{\bf k}}}\braket{u_{m{\bf k}}}{\partial_{k_{j}}u_{p{\bf k}}}}$ over momentum space, with its real part, commonly known as the non-Abelian quantum metric, being $[\textbf{g}^{(m)}_{ij}]_{np}=\text{Re}~{}[\textbf{Q}^{(m)}_{ij}]_{np}$ Provost and Vallee (1980); Ma et al. (2010); Bouhon et al. (2023). The multiband quantum metric encodes the optical transition dipole matrix elements, $[\textbf{g}^{(m)}_{xx}]_{nn}=|\bra{\psi_{n{\bf k}}}\hat{x}\ket{\psi_{m{\bf k}}}|^{2}$ Ahn et al. (2020, 2021), which is at the heart of both linear and nonlinear responses. The multiband quantum metric of electrons realizes lower bounds due to 1D crystalline topologies that translate into minimal values of (higher-order) optical susceptibilities,
$\langle g_{xx}\rangle\equiv\sum^{\text{unocc}}_{m}\sum_{k}\text{Tr}~{}{\textbf{g}}^{(m)}_{xx}\geq\frac{a^{2}}{4}\nu_{\text{FK}},$ (1)
where $\langle g_{xx}\rangle$ is band-averaged quantum metric and ${\nu_{\text{FK}}\in\mathbb{Z}_{2}}$ the Fu-Kane parity invariant protected by the $\mathcal{I}$-symmetry Fu and Kane (2007). The bound can be derived from (i) non-trivial inversion symmetry indicators pinpointing how the irreducible representations of the obstructed bands transform Fu and Kane (2007); Kruthoff et al. (2017); Po et al. (2017); Bradlyn et al. (2017), (ii) translating these into real-space projectors Herzog-Arbeitman et al. (2022), and (iii) using an optimization argument of Ref. Herzog-Arbeitman et al. (2022). The technical details of the derivation are given in Sec. I of the Supplemental Material (SM) SI . In the context of the maximally localized Wannier functions, $\langle g_{xx}\rangle=\langle{\Delta x^{2}}\rangle$ Marzari and Vanderbilt (1997). Hence, the invariant imposes a lower-bound constraint on the spread of the wavefunction in the real-space basis. In topological crystalline phases, any topologically nontrivial band realizes $\nu_{\text{FK}}=1$. In the following, we demonstrate that this bound yields a topological enhancement of the multiband geometry-dependent response functions within the first and third-order optical responses.
First- and third-order optical responses.— The linear optical conductivity of a band insulator in the first-order response to an electric field of frequency $\omega$ reads Ahn et al. (2021),
$\sigma_{xx}(\omega)=\frac{e^{2}}{2\hbar^{2}}\int_{\text{BZ}}\frac{\mathop{}\!\mathrm{d}k}{2\pi}\sum_{n,m}~{}f_{nm}E_{mn}\text{Tr}~{}{\textbf{g}}^{(m)}_{xx}\delta(\omega-E_{mn}/\hbar),$ (2)
where $f_{nm}=f_{n}-f_{m}$ are the thermal Fermi-Dirac occupation factors and $E_{mn}$ is the difference in energy between valence and conduction bands $n$ and $m$, respectively. For crystals possessing the same band structure, the distinct first order optical conductivity is purely determined by the multiband quantum metric. This allows one to uniquely distinguish band topologies by inducing the corresponding band geometries. Importantly, changes in band topology, as realized through external parameters such as lattice strain, as shown in Fig. 1(a,b), result in changes in the optical response. As we shall discuss in the representative case of $\pi$-conjugate chains and related one-dimensional crystals, lattice strain can be leveraged to engineer the opacity and optical nonlinearities, as illustrated in Fig. 1(c).
While the second-order optical responses vanish under the $\mathcal{I}$-symmetry in centrosymmetric crystals, third-order responses remain finite and can be controlled with band geometry. In the static limit, the third-order susceptibility can be enhanced quadratically with the considered multiband quantum metric
$\chi^{(3)}=\frac{e^{4}}{\epsilon_{0}}\int_{\text{BZ}}\frac{\mathop{}\!\mathrm{d}k}{2\pi}~{}\sum_{n,m,p,q}\frac{[\textbf{g}^{(p)}_{xx}]_{nm}[\textbf{g}^{(q)}_{xx}]_{mn}}{E_{mn}E_{pn}E_{qn}},$ (3)
where $\epsilon_{0}$ is the vacuum permittivity. The derivation of Eq. (3) is provided in Sec. II of the SM SI .
While the enhancement of $\chi^{(3)}$ can be induced by nontrivial quantum geometry in the numerator, it can be further achieved through the tunable band gap magnitudes in the denumerator. Notably, materials with identical band gaps, yet with distinct crystalline topologies, realize distinguishable numerators, with topological phase experiencing a quadratic enhancement. By combining this enhancement with Eq. (1) and sum rules, we derive a fundamental topological lower bound on the hyperpolarizability of a nontrivial insulator,
$\chi^{(3)}\geq\alpha^{(3)}\times\frac{a^{10}}{n^{3}},$ (4)
where $\alpha^{(3)}=\frac{1}{2}\pi^{3}e^{4}m_{e}^{3}\epsilon^{-1}_{0}\hbar^{-6}=6\times 10^{50}~{}\text{m}^{-5}/\text{V}^{2}$, $m_{e}$ is the electron mass, and $n$ is the number density of electrons in the crystal. The derivation of Eq. (4) employs (i) the energy gap bound of Ref. Kivelson (1982) due to quantum metric (ii) combined with the sum rules of Ref. Souza et al. (2000), as employed in the linear susceptibility bound of Ref. Onishi and Fu (2024). The technical details of the derivation are given in Sec. IV of the SM SI .
From models to real materials.— We next employ an effective, two-band Hamiltonian to capture the geometry induced by $\mathbb{Z}_{2}$ obstructed topology in the presence of lattice strain. This model is schematically depicted in Fig. 1(a). The Hamiltonian that minimally realizes such general phenomenology on a quasi-one-dimensional chain takes the form of a strained SSH model Su et al. (1979, 1980),
$H(k_{x},\lambda)=\begin{pmatrix}0&t_{1}(\lambda)+t_{2}(\lambda)e^{-ik_{x}a(\lambda)}\\ t_{1}(\lambda)+t_{2}(\lambda)e^{ik_{x}a(\lambda)}&0\\ \end{pmatrix},$ (5)
where $\lambda>0$ is an external tensile stress parameter, ${a(\lambda)\equiv a_{1}(\lambda)+a_{2}(\lambda)}$ is the strain-renormalized lattice parameter with bond lengths $a_{j}(\lambda)=a_{j}(0)[1+\lambda]$, and ${t_{j}(\lambda)=t_{j}(0)e^{-\gamma_{j}[a_{j}(\lambda)-a_{j}(0)]}}$ are the strain-dependent hopping amplitudes, where $\gamma_{j}$’s parameterize bond stiffness Tepliakov et al. (2023). We validate our strain-dependent models with hopping parameters obtained from first-principles calculations Jankowski et al. (2025). We correspondingly present our numerical results in Fig. 2. In Fig. 2(a), we show how the profile of quantum metric $g_{xx}$ over $k$-space changes with the strain $\lambda$. The metric supporting the topological lower bound is enhanced at the band edge $k=\pi/a$, as the dominant contributions arise at the band inversion region, where the local derivatives of the Bloch eigenvectors winding across the BZ are maximized. Upon crossing the critical strain $\lambda=\lambda_{c}$, the band edge metric contributions provides for an average $\langle g_{xx}\rangle$ that satisfies the bound in the topological phase ($\nu_{\text{FK}}=1$). Fig. 2(b) presents a dependence of the optical weight, i.e., weighted integrated optical conductivity, or equivalently, integrated interband contribution to the imaginary part of dielectric constant ${W_{xx}=\int^{\infty}_{0}~{}\mathop{}\!\mathrm{d}\omega~{}\frac{\sigma_{xx}(\omega)}{\omega}=\frac{\pi e^{2}}{\hbar}\langle g_{xx}\rangle}$, which is also lower-bounded by the topological invariant. In. Fig. 2(c), we benchmark the first-order optical conductivity and show how its contributions to the bounded optical weights gradually arise. Fig. 2(d) shows how the colossal values of $\chi^{(3)}$ arise with strain $\lambda$, before vanishing sharply upon transitioning to the trivial phase across the critical point. We remark that the value of $\chi^{(3)}$ at vanishing $\lambda$ is in excellent agreement with experimental observations. Importantly, the bound of the quantum metric, as manifested in Fig. 2(c-d), provides for the robustness of lower-bounded optical responses of topological phases, especially away from the singular critical points that correspond to the strain-induced phase transitions. Further details on the model-dependent parameterization and controllability of quantum geometry within $\pi$-conjugated chains are given in Sec. III of SM SI .
Figure 3:  Examples of materials for crystalline topology-induced, quantum-geometrically enhanced nonlinear optics. From top to bottom: polyacetylene, poly-($p$-phenylene), polypentacene, and graphene nanoribbons. Under external control, such as lattice strains, these materials can realize $\mathcal{I}$-symmetry protected nontrivial $\mathbb{Z}_{2}$ topology and topologically-enhanced hyperpolarizability $\chi^{(3)}$, which culminates in nonlinear polarization $P^{(3)}$. Further chemical functionalizations of such $\mathbb{Z}_{2}$-topological materials, e.g., through perhalogenations, can provide a further knob to tune the quantum metric and higher-order susceptibilities.
We stress that this model Hamiltonian and accompanying results are effectively realized by the low-energy bands of an ample variety of one-dimensional $\pi$-conjugate systems, such as polyacetylene Su et al. (1979, 1980), polyacenes Cirera et al. (2020), polyphenylenes, and graphene nanoribbons Tepliakov et al. (2023), as overviewed in Fig. 3. Their low-energy valence and conduction bands are universally governed by an effective physics of hopping between the Wannier orbitals $A,B$ that are predominantly constituted by the weighted combinations of the atomic orbitals with $p_{z}$ character Su et al. (1979); Cirera et al. (2020); Jankowski et al. (2025). The physical implication of our findings is that the materials shown in Fig. 3 are expected to exhibit quantum-geometrically enhanced hyperpolarizabilities that originate from their nontrivial, inversion symmetry-protected band topology. Indeed, the introduction of additional bands does not alter out conclusions, as we demonstrate in Sec. V of the SM SI , where we also provide an extended discussion of multiband models. The reason for this traces back to the fact that the additional bands will experience higher energy gaps $E_{mn}$, suppressing the $\chi^{(3)}$ contributions as $1/E^{3}_{mn}$. Furthermore, the metric similarly scales with $1/E^{2}_{mn}$ [cf. Fig. 2(a)], which for topologically trivial additional bands provides only a marginal contribution to the numerator of $\chi^{(3)}$.
From quantum to semiclassical picture.— We now present an intuitive picture of the topologically-induced quantum geometric phenomenology. Semiclassically, electrons in band gap insulators can be modeled with an effective nonlinear anharmonic oscillator Hamiltonian, ${H=\frac{p^{2}}{2m_{e}}+U(x)}$ with a $\mathcal{I}$-symmetric confining nonlinear potential ${U(x)=\frac{1}{2}\kappa x^{2}+\frac{1}{4}bx^{4}}$, where $\kappa$ is an effective spring constant, $\kappa=m_{e}\omega_{0}^{2}$, in which $\omega_{0}$ is a resonant frequency and $b$ quantifies the degree of anharmonicity. Notably, ${\kappa~{}\propto 1/\langle g_{xx}\rangle=1/\langle\Delta x^{2}\rangle}$ Onishi and Fu (2024), where $\langle\Delta x^{2}\rangle$ is a spatial spread of a localized state wavefunction, as representable in terms of Wannier functions Marzari and Vanderbilt (1997). From the theory of nonlinear anharmonic oscillator discussed in Sec. VI of the SM SI , one expects $\chi^{(3)}\propto b/\kappa^{4}$, by recognizing that, typically, $b\propto\kappa/\langle g_{xx}\rangle$, explaining the effective scaling of the ${\chi^{(3)}\propto\langle g_{xx}\rangle^{2}}$ within a full quantum-mechanical treatment. The semiclassical picture, in connection to quantum geometry, further supports the interpretation of our key result. Namely, the topological symmetry-indicated winding of the Bloch states in a one-dimensional insulator: (i) induces a geometric spread in the localization of the bounded electrons, which (ii) semiclassically reflects an effective spring constant and its inteplay with anharmonicity. The latter determine the nonlinear optical response of the driven localized electrons within an anharmonic oscillator picture.
Discussion and conclusions.— In summary, we have shown that the colossal linear and higher-order optical responses observed in one-dimensional $\pi$-conjugate chains can be understood by means of the modern notions of band topology and quantum geometry, where the latter is bounded from below by the former. We have established a transparent connection of the multiband quantum geometry with the third-order optical responses, demonstrating that the crystalline-symmetry-protected topology induces a lower bound on the hyperpolarizability. Our quantum-geometric perspective sheds a new light on the long-standing controversy of thediscrepancy betweeen $\chi^{(3)}$ of the cis and trans geometric isomers of polyacetylene  Sinclair et al. (1988); Yu et al. (1989b); Bredas et al. (1994). The distinction between cis- and trans-polyacetylene arises from the different spatial arrangements of atoms with identical chemical connectivity, which induces inequivalent band topology in the two isomers. The band topology classified by the winding of the Bloch eigenvectors provides for a lower bound in quantum geometric tensor and in the associated enhanced nonlinear optical response matrix elements in the trans isomer. While being manifestly larger in the topological polymer, the metric band average lies above the topological lower bound and does not saturate, unless the band dispersion is forced to approach the flat band limit. The minimization of lower-bounded quantum metric in such limiting case is consistent with the other known flatness conditions for the metric bound saturations Roy (2014); Kwon and Yang (2024); Wahl et al. (2025). In the provided picture, cis-polyacetylene is topologically trivial (${\nu_{\text{FK}}=0}$), hence, realizes no bound on geometry with average quantum metric $\langle g_{xx}\rangle$ that induces an enhancement of the $\chi^{(3)}$, unlike in the topological trans-polyacetylene ($\nu_{\text{FK}}=1$).
The theory presented in this work offers a step forward by revealing that the origin of the $p_{z}$-orbital band polarizability is topologically-induced, as associated with the lower-bounded quantum metric providing for an enhanced spread of the electronic wavefunction in the topological phase. To the best of our knowledge, the importance of band topology and geometry for nonlinear responses in the static limits— especially concerning hyperpolarizabilities—-were not addressed in previous studies. Given that any band topology is fully encoded in the non-Abelian Berry connections, which are fully captured by the quantum geometry determining nonlinear optical responses Ahn et al. (2021), our conceptual framework consistently and comprehensively captures the role of band topology for hyperpolarizabilities in any class of materials. In addition to offering an understanding of the observed difference in the hyperpolarizabilities between distinct isomers through the notion of quantum geometric bound, the phenomenological picture underlying our findings translates into topologically-induced quantum lattice fluctuations Yu et al. (1989a); Takahashi (1995, 1997b) in the form of virtual soliton-antisoliton pairs Yu et al. (1989a). Following identifications of positional charge correlators with electronic localization lengths and quantum geometry of Refs. Kivelson (1982); Souza et al. (2000); Tam et al. (2024), such quantum charge density fluctuations are driven by quantum metric.
To conclude, our work demonstrates that topologically lower-bounded higher-order geometric optical responses can be achieved in materials possessing nontrivial crystalline topology, thus opening a pathway to identify and engineer novel platforms for enhanced nonlinear phenomena within the growing family of topologically classified materials.
I Acknowledgements
Acknowledgements.
W.J.J. and R.-J.S. acknowledge Bartomeu Monserrat and Joshua J.P. Thompson for stimulating discussions. W.J.J. acknowledges funding from the Rod Smallwood Studentship at Trinity College, Cambridge. R.-J.S. acknowledges funding from an EPSRC ERC underwrite grant EP/X025829/1 and a Royal Society exchange grant IES/R1/221060 as well as Trinity College, Cambridge.